Fermi National Laboratory


Accelerator Update

Friday April 2
The day shift began with Tevatron (TeV) experts investigating a quench, with the Antiproton Source (Pbar) stacking, with the Main Injector (MI) and Recycler conducting studies, and with beam to MiniBooNE.

Operations established beam to the Meson MTest experiment at 8:40 AM.

Operators prepared the Meson MCenter experiment for an access at 9:38 AM.

At 9:46 AM, operators prepared D-zero (D0) collision hall for an access.

Cryo system technicians reported at 10:32 AM that there was a leaking Kautzky valve at TeV sector D11. TeV sector B11 also shows the same indication. Operators made the TeV ready for access to investigate those problems.

Operators prepared the Collider Detector at Fermilab (CDF) for an access at 11:52 AM. They came out of access thirty-minutes later.

At 3:10 PM, Operations reestablished beam to MCenter.

Operators reported at 3:17 PM that they had discovered the cause for the TeV quench. It was due to a TeV sector D1 bulk power supply not operating properly. Operators contacted a power supply tech to make repairs.

Operations put the TeV through a dry squeeze at 3:46 PM.

TeV experts began a study period at 4 PM.

Operations turned off the MTest experiment at 5:13 PM.

Operations put the TeV through a wet squeeze at 10:57 AM. TeV experts finished their studies at 11:20 PM and Operations put the TeV into shot setup at 11:35 PM.

Saturday April 3
The midnight shift began with the TeV in shot setup, with the beam off to MCenter until after the TeV shot, and with beam to MiniBooNE.

At 12:34 AM, a Booster kicker (MKS04) tripped off. An expert came in at 1:31 AM, found the problem, and had it repaired within twenty minutes.

Operations established store 3348 at 3:46 AM with an initial luminosity of 50.85E30. Pbar resumed stacking.

At 3:54 AM, Operations reestablished beam to Switchyard (SY) and to MCenter.

A MI expert began a slow acceleration study at 1:25 PM.

The LCW pumps at MI-52 tripped off at 5:50 PM. Operators investigated and managed to get the system back on within thirty minutes. Pbar stacking resumed at 6:12 PM.

The MI study ended at 8:47 PM.

Sunday April 4
The midnight shift began with Operations monitoring store 3348, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 4:40 AM, Operations reported having some controls issues due to the time change. Operators called an expert who rebooted some systems.

Operations terminated store 3348 at 7:41 AM.

Operations put the TeV into shot setup at 8:34 AM.

At 10:54 AM, after all the protons and antiprotons had been loaded into the TeV for the next store, the TeV suffered a sector A1 quench. A Pbar abort kicker (AAK2) prefired and caused the quench. Pbar resumed stacking.

Cryo system techs at CHL reported at noon that they were hit by a power glitch. Operations asked the Duty Electrician to investigate. A Meson water system tripped off at the same time. Operators called in a tech to check out the water system before turning it back on.

At 1:58 PM, the High Voltage expert reported that feeder 35 had faulted. Power was being moved to feeder 32.

CHL reported at 2:19 PM that they had recovered from the power glitch.

Operations restored beam to MCenter at 3:56 PM.

TeV experts began a study period at 7:17 PM.

Operators noticed at 10:10 PM that the TeV sector A3 wet engine wasn't regulating properly. Upon investigation operators saw that the flywheel wasn't rotating properly. This halted the TeV shot setup until experts could arrive and make repairs.

Monday April 5
The midnight shift began with the TeV shot setup halted due to wet engine problems, with Pbar stacking, with beam to MiniBooNE, and with SY conducting studies.

At 12:32 AM, cryo system techs reported that they needed to replace the flywheel and belt.

At 1:56 AM, operators investigated a problem with the Linac 400 MeV Debuncher (Booster) water regulation. The operators discovered that the problem was with the control program and so they set the valve manually to maintain the nominal 90 F temperature.

Cryo system techs had the A3 flywheel replaced by 2:11 AM.

Operation put the TeV into shot setup at 6:18 AM.

Today's Plans The plans for today, after establishing the next store, include a MI access to test a NuMI magnet and power supply, and the stashing of antiprotons to the Recycler.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/5/2004   email Fermilab